OpenStudy (anonymous):

solve the limit; lim x'2 -4/x'3 -8 x→-2

myininaya (myininaya):

x'?

OpenStudy (anonymous):

X^

OpenStudy (anonymous):

power

myininaya (myininaya):

oh like x^2 x^3

OpenStudy (anonymous):

yeah, i dont know how to put that with the keyboard

myininaya (myininaya):

\[\lim_{x \rightarrow -2}\frac{x^2-4}{x^3-8}=\lim_{x \rightarrow -2}\frac{(x-2)(x+2)}{x^3-8}=\lim_{x \rightarrow -2}\frac{(x-2)(x+2)}{(x-2)(x^2+2x+4)}\]

myininaya (myininaya):

the x-2's cancel

OpenStudy (anonymous):

i have x+2/x'2+2x+4, but i dont know what to do next

myininaya (myininaya):

\[\lim_{x \rightarrow -2}\frac{x+2}{x^2+2x+4}=\frac{-2+2}{(-2)^2+2(-2)+4}=0\]

OpenStudy (anonymous):

yes that's it, but is indeterminate

OpenStudy (anonymous):

how is it indeterminate

myininaya (myininaya):

no

myininaya (myininaya):

we dont have 0/0

OpenStudy (anonymous):

he canceled out

myininaya (myininaya):

she*

myininaya (myininaya):

we have 0/(4-4+4)=0/4=0

OpenStudy (anonymous):

oh i see, then is correct, he*

OpenStudy (anonymous):

she* canceled out stuff so that it was no longer indeterminate form when youget an indeterminate form in calculus 1, the only way ou can solve it is by manipulation of the problem

myininaya (myininaya):

we had 0/0 at the begining but like out said i canceled the x-2's so no more 0/0

OpenStudy (anonymous):

whether that be cancelling factors or rationalizing

myininaya (myininaya):

also do you know l'hospital if you know way i can show you another way to find the limit here

OpenStudy (anonymous):

ok, thanks, i was confused

OpenStudy (anonymous):

please, show me how

myininaya (myininaya):

oops wait don't look at that

OpenStudy (anonymous):

oh, a derivative

myininaya (myininaya):

wiat what's happening what i'm doing

myininaya (myininaya):

didn't i get 0 above?

OpenStudy (anonymous):

is 0/12?

myininaya (myininaya):

\[\lim_{x \rightarrow -2}\frac{x^2-4}{x^3-8}=\lim_{x \rightarrow -2}\frac{2x}{3x^2}=\lim_{x \rightarrow -2}\frac{2}{3x}=\frac{2}{3(-2)}=\frac{1}{-3}\] ok something is wrong with one of my ways

myininaya (myininaya):

let me look above

myininaya (myininaya):

omg out what's wrong with what i did?

myininaya (myininaya):

out are you still there?

myininaya (myininaya):

let me get zarkon

OpenStudy (anonymous):

i don't get it, but you were right at the top

myininaya (myininaya):

ok i did it two ways and i didn't get the samething both ways so either im confused about calculus or algebra

OpenStudy (amistre64):

(-2)^3 = -8 for starters

myininaya (myininaya):

omg no way

OpenStudy (zarkon):

you can't use L'Hospitals unless you have an indeterminant form

myininaya (myininaya):

i suck at algebra

OpenStudy (zarkon):

look like we do

myininaya (myininaya):

i thought i had 0/0 mu we didn't have 0/0 lol

OpenStudy (zarkon):

except the second time you use l'hospitals rule

myininaya (myininaya):

so we can't use l'hospital way sorry

myininaya (myininaya):

i used l'hospital once

OpenStudy (amistre64):

lHopital is effective on indeterminant forms :)

OpenStudy (zarkon):

lol..nope you don't

OpenStudy (amistre64):

\[\lim_{x->-2}\frac{x^2-4}{x^3-8}=\frac{(-2)^2-4}{(-2)^3-8}=\frac{0}{-16}=0\]

OpenStudy (anonymous):

at the beginning